# MObI: Multimodal Object Inpainting Using Diffusion Models

### v1

@@ Abstract
Safety-critical applications, such as autonomous driving, require extensive multimodal data for rigorous testing. Methods based on synthetic data are gaining prominence due to the cost and complexity of gathering real-world data but require a high degree of realism and controllability in order to be useful. This paper introduces MObI, a novel framework for Multimodal Object Inpainting that leverages a diffusion model to create realistic and controllable object inpaintings across perceptual modalities, demonstrated for both camera and lidar simultaneously. Using a single reference RGB image, MObI enables objects to be seamlessly inserted into existing multimodal scenes at a 3D location specified by a bounding box, while maintaining semantic consistency and multimodal coherence. Unlike traditional inpainting methods that rely solely on edit masks, our 3D bounding box conditioning gives objects accurate spatial positioning and realistic scaling. As a result, our approach can be used to insert novel objects flexibly into multimodal scenes, providing significant advantages for testing perception models. Code will be made available upon publication.
@@ 1 Introduction
For example, reference-based image inpainting methods [64, 5, 45, 24] can produce realistic samples that seamlessly blend into the scene using a single reference, but they often lack precise control over the 3D positioning and orientation of the inserted objects. In contrast, methods based on actor insertion using 3D assets [52, 4, 73, 55, 6, 29, 11, 25] provide a high degree of control—enabling precise object placement in the scene—but often struggle to achieve realistic blending and require high-quality 3D assets, which can be challenging to produce. Similarly, reconstruction methods [54, 50, 66] are also highly controllable but require almost full coverage of the inserted actor. We illustrate some of these shortcomings in Fig. 1. While more recent methods have explored 3D geometric control for image editing [53, 59, 67, 39, 35, 68], none consider multimodal generation which is crucial in autonomous driving. We provide an extended overview of the prior art in Sec. A.
@@ 2.2 Multimodal generation
We introduce a modality-dependent cross-modal attention which looks at the tokens of the other modality from the same scene in the batch. We derive the query, key, and value representations from the input camera and lidar features with layer normalisation applied for cross-attention from camera to lidar. Using learnable transformations $W_{Q}^{\text{(C)}},W_{K}^{\text{(R)}},W_{V}^{\text{(R)}}$, we compute the cross-attention as:
$\text{Attn}^{\text{(C)}}=\text{softmax}\left(\frac{Q^{\text{(C)}}(K^{\text{(R)}})^{T}}{\sqrt{d_{\text{head}}}}\right)V^{\text{(R)}},$
where $Q^{\text{(C)}}=W_{Q}^{\text{(C)}}\mathbf{h}^{\text{(C)}}$, $K^{\text{(R)}}=W_{K}^{\text{(R)}}\mathbf{h}^{\text{(R)}}$, and $V^{\text{(R)}}=W_{V}^{\text{(R)}}\mathbf{h}^{\text{(R)}}$. We then update the camera features by adding a residual connection through a zero-initialised gating module: $\mathbf{h}^{\text{(C)}}\leftarrow\mathbf{h}^{\text{(C)}}+\text{Gate}^{\text{(C)}}(\text{Attn}^{\text{(C)}}).$ The computation for lidar-to-camera cross-attention is analogous, with lidar features attending to the camera modality. We do not restrict the cross-modal attention and let the network learn an implicit correspondence which is facilitated by the respective projected bounding boxes. Lastly, we concatenate the camera and lidar tokens within the batch.
@@ 2.3 Inference and compositing
Final results are obtained by compositing the edited camera and range images back into the original scene. For images, we extract the region within the projected bounding box from the edited image $\tilde{\mathbf{x}}^{(\text{C})}$ and insert it back into the destination image $I_{d}$. Following the approach of POC [7], a Gaussian kernel is applied to improve blending, resulting in the final composited image. For lidar, we create a 2D mask $\mathbf{m}_{\text{points}}$ by selecting points from the original lidar point cloud $P_{d}$ that fall within the destination 3D bounding box. The edited range image $\mathbf{\tilde{x}}^{\text{(R)}}$ is then resized to an object-centric range view using average pooling and denormalised before computing coordinate and intensity values (see Sec. B.2). We then replace pixels in the original range view $R_{d}$ with the corresponding pixels from the edited range image if either (i) they fall within $\mathbf{m}_{\text{points}}$ or (ii) its corresponding 3D point in the edited range image is contained by the bounding box of the object.
@@ 2.4 Training details
We consider objects from the nuScenes dataset [3] train split with at least 64 lidar points, whose 2D bounding box is at least $100\times 100$ pixels, with a 2D IoU overlap not exceeding 50%, and current camera visibility of at least 70%. Unless stated otherwise, our model is trained on “car” and “pedestrian” categories, dynamically sampling 4096 new actors per class each epoch. During training, once an object is selected, its current scene serves as the destination, from which we extract the 3D bounding box, environmental context, and ground truth insertion.
@@
Object references are taken from the same object at a different timestamp picked randomly as follows. We collect references for the current object across all frames that meet the previous criteria to ensure good visibility and arrange them by normalised temporal distance $\Delta t$, where $1$ represents the furthest reference in time and 00 represents the current one. We then sample references randomly using a beta distribution $\Delta t\sim\text{Beta}(4,1)$ which ensures a preference for instances of the object that are far away from the current timestamp, see Sec. B for details.
@@ 3.1 Object insertion and replacement
In order to avoid situations where inpainted objects are placed at locations incompatible with the scene (e.g. a car on the pavement), we use the position of existing objects and perform either object reinsertion or replacement, which differ by the choice of the inpainting reference. Doing so, we test the model’s ability to generate realistic objects conditioned on a 3D bounding box while being semantically consistent with the scene. We sample 200 original objects from the nuScenes val set as in Sec. 2.4, balanced across the “car” and “pedestrian” classes.
@@
We define two types of references: same reference, where the source and destination images and bounding boxes are identical, meaning the object is reinserted in the exact same scene and position; and tracked reference, where the object is reinserted given its reference from a different timestamp, using the same sampling strategy described in Sec. 2.4. This setting tests if the model can preserve the object’s appearance, and realistically perform novel view synthesis (for tracked reference).
@@
Results are presented on Fig. 3 both for replacement (rows 1–4) and insertion (row 5). We see that inpainted objects correspond tightly to their conditioning 3D bounding boxes while having a high degree of realism, both for camera (RGB) and lidar (depth and intensity), and show a strong coherence (lightning, weather conditions, occlusions, etc.) with the rest of the scene. The last row showcases object deletion, which can be achieved by using an empty reference image (note that we use empty references during training, as described in Sec. 2.4). Even though references in the replacement setting are from a different domain (time of day/weather) the model is able to inptaint such objects realistically. See Fig. S3 for more examples, including failure cases. Finally, we illustrate the flexibility of our bounding box conditioning and show that it is able to generate multiple views with a high degree of consistency, as illustrated on Fig. 4, Fig. S1 and Fig. S5.
@@
Figure 4: Our method can generate multiple novel views from a single reference image while maintaining multimodal consistency. From left to right: reference image $\mathbf{x}_{\text{ref}}$, extracted from a separate source scene; original destination scene with the original RGB image $\mathbf{x}^{\text{(C)}}$ and lidar range depth $\mathbf{x}_{0}^{\text{(R)}}$; and edited scenes. For extended results including intensity, check Fig. S1.
@@
Table 1: Adaptation methods of the pre-trained image VAE [21] from StableDiffusion [42] showing improved lidar reconstruction for depth and intensity. Depth is reported in meters and intensity is on a scale of $[0,255]$.
@@ Results
We report all realism metrics for camera-lidar object inpainting in Tab. 2 for the reinsertion and replacement settings. Compared to camera-only inpainting methods, MObI ($D=512$) achieves better results than PbE [64] across almost all benchmarks. We note that this method achieves competitive results in terms of FID, producing samples which are close in distribution to the target ones, yet LPIPS is much worse. This perceptual misalignment, which is more severe than even MObI with $D=256$ with no bounding box conditioning, might indicate that the use of joint generation of camera and lidar improves semantic consistency within the scene. We also compare to simple copy&paste and show that this produces unrealistic composited images when replacing objects, even though it is sometimes used for training object detectors [12, 8, 51, 72]. Note that object reinsertion results for copy&paste, as well as CLIP-I scores, are not computed as the comparison would be unfair in this case. We provide additional realism metrics in Tab. S1
We provide ablations for the 3D bounding box and the gated cross-attention adapter for $D=256$. Without the adapter, the box token is concatenated with the reference token in the PbE [64] cross-attention layer, followed by direct fine-tuning. Due to the lack of a baseline for lidar object inpainting realism, we provide comparative results for all experiments and ablation in the hope that this could constitute a baseline for future work. Comparing MObI (256) with both bbox conditioning and adapter to the one without bbox conditioning we notice significant improvements in perceptual alignment. The use of gated cross-attention leads to more realistic samples in the camera space, yet it does not improve lidar, hinting towards differences in training regimes for the two modalities. Finally we note that realism scales strongly with resolution, leading us to believe that models operating at a larger resolution would improve realism even more.
@@
Figure 5: Detection performance of an off-the-shelf BEVFusion [33] object detector on objects reinserted using our method. Left: we compute mAP at the scene-level, and TP errors (translation, scale, and orientation) on the reinserted objects only. Left: distribution of the scores of the true-positives show a modest shift towards lower scores for edited objects.
@@ 3.3 Object detection on reinserted objects
We compute mAP and error metrics on the re-inserted objects. Scene-level metrics such as mAP can not be easily restricted to edited objects [^2] This is because such metrics usually require false-positives which are detections that have not been matched to any ground truth-object within a scene, but not to a specific subset of ground-truth objects. and will not be very sensitive to detection errors on these objects. This is why we complement mAP with true-positive error metrics restricted to the re-inserted objects, which are computed following the usual matching procedure from the nuScenes devkit [3] but considering only the ground-truth/detection pairs that correspond to inpainted objects.
@@ Results
Object detection results are presented on Fig. 5 (left), and we see that reinsertion comes at a small cost in object detection performance but that errors remain small (e.g. 0.161 AOE corresponds to a $9^{\circ}$ average error) while scene-level mAP is very similar. We also show the distribution of the scores of the true-positives on Fig. 5 (right), where we can see that the scores suffer a modest decrease when the detector is applied to the reinserted samples. Overall, this shows that while a small domain gap exists, our bounding box conditioning is able to produce samples that are both realistic and accurate, and that off-the-shelf detector can successfully detect such objects even though they have not been trained on any synthetic data generated by our method. We show a sample of detections in Fig. S2 where the reinserted object is detected accurately and the bounding boxes of the untouched objects remain almost identical.
@@ 4 Strengths, limitations, and future work
Because we leverage a diffusion model pre-trained on webscale data, combined with the CLIP image encoding to guide the appearance of the inpainted object, our model can generate objects for classes unseen during training. In Fig. S4, we replace objects from nuScenes [3] classes beyond “car” and “pedestrian”, producing plausible, yet seemingly lower-quality results compared to classes seen during training.
Inpainting completely out-of-domain references produces unpredictable results. Since the model is finetuned on a narrow domain, the model reverts to known classes when faced with unfamiliar references, such as turning a horse into a brown car in Fig. S4333It is sadly impossible for us to tell if this car has a lot of horsepower.. Extending our method to a true open-world setting by training on multiple 3D object detection datasets, as in [36], is an exciting future direction.
@@ A Extended Related Work
Synthetic lidar data generation has also advanced significantly. LidarGen [74] and LiDM [41] employ diffusion for lidar generation, with the latter also incorporating semantic maps, bounding boxes, and text. UltraLidar [62] densifies sparse lidar point clouds, while RangeLDM [19] accelerates lidar data generation by converting point clouds into range images using Hough sampling and enhancing reconstruction through a range-guided discriminator. DynamicCity [2] generates lidar sequences conditioned on dynamic scene layouts, and [60] generates object-level lidar data, demonstrating its benefits for object detection. However, these works do not jointly generate camera and lidar data, and full-scene generation can result in a large domain gap, particularly for downstream tasks like object detection, making it challenging to create realistic counterfactuals.
@@ B.3 Additional training details
Ground Truth Vanilla Ours
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Original Reference Insertion (C) Insertion (R) - depth Insertion (R) - intensity
Original Reference Insertion (C)
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Original Reference Replacement (C)
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption

### v2

@@ Abstract
Safety-critical applications, such as autonomous driving, require extensive multimodal data for rigorous testing. Methods based on synthetic data are gaining prominence due to the cost and complexity of gathering real-world data but require a high degree of realism and controllability in order to be useful. This paper introduces MObI, a novel framework for Multimodal Object Inpainting that leverages a diffusion model to create realistic and controllable object inpaintings across perceptual modalities, demonstrated for both camera and lidar simultaneously. Using a single reference RGB image, MObI enables objects to be seamlessly inserted into existing multimodal scenes at a 3D location specified by a bounding box, while maintaining semantic consistency and multimodal coherence. Unlike traditional inpainting methods that rely solely on edit masks, our 3D bounding box conditioning gives objects accurate spatial positioning and realistic scaling. As a result, our approach can be used to insert novel objects flexibly into multimodal scenes, providing significant advantages for testing perception models. Project page: https://alexbubu.com/mobi
@@ 1 Introduction
For example, reference-based image inpainting methods [65, 5, 46, 25] can produce realistic samples that seamlessly blend into the scene using a single reference, but they often lack precise control over the 3D positioning and orientation of the inserted objects. In contrast, methods based on actor insertion using 3D assets [53, 4, 74, 56, 6, 30, 11, 26] provide a high degree of control—enabling precise object placement in the scene—but often struggle to achieve realistic blending and require high-quality 3D assets, which can be challenging to produce. Similarly, reconstruction methods [55, 51, 67] are also highly controllable but require almost full coverage of the inserted actor. We illustrate some of these shortcomings in Fig. 1. More recent methods have explored 3D geometric control for image editing [54, 60, 68, 40, 36, 69], as well as object-level lidar generation [22]. However, none consider multimodal generation, which is crucial in autonomous driving. We provide an extended overview of the prior art in Sec. A.
@@ 2.2 Multimodal generation
We introduce a modality-dependent cross-modal attention which looks at the tokens of the other modality from the same scene in the batch. We derive the query, key, and value representations from the input camera and lidar features with layer normalisation applied for cross-attention from camera to lidar. Using learnable transformations $W_{Q}^{\text{(C)}},W_{K}^{\text{(R)}},W_{V}^{\text{(R)}}$, we compute the cross-attention as: $\text{Attn}^{\text{(C)}}=\text{softmax}\left({Q^{\text{(C)}}(K^{\text{(R)}})^{T}}/{\sqrt{d_{\text{head}}}}\right)V^{\text{(R)}},$ where $Q^{\text{(C)}}=W_{Q}^{\text{(C)}}\mathbf{h}^{\text{(C)}}$, $K^{\text{(R)}}=W_{K}^{\text{(R)}}\mathbf{h}^{\text{(R)}}$, and $V^{\text{(R)}}=W_{V}^{\text{(R)}}\mathbf{h}^{\text{(R)}}$. We then update the camera features by adding a residual connection through a zero-initialised gating module: $\mathbf{h}^{\text{(C)}}\leftarrow\mathbf{h}^{\text{(C)}}+\text{Gate}^{\text{(C)}}(\text{Attn}^{\text{(C)}}).$ The computation for lidar-to-camera cross-attention is analogous, with lidar features attending to the camera modality. We do not restrict the cross-modal attention and let the network learn an implicit correspondence which is facilitated by the respective projected bounding boxes. Lastly, we concatenate the camera and lidar tokens within the batch.
@@ 2.3 Inference and compositing
Final results are obtained by compositing the edited camera and range images back into the original scene. For images, we extract the region within the projected bounding box from the edited image $\tilde{\mathbf{x}}^{(\text{C})}$ and insert it back into the destination image $I_{d}$. Following the approach of POC [7], a Gaussian kernel is applied to improve blending, resulting in the final composited image. For lidar, we create a 2D mask $\mathbf{m}_{\text{points}}$ by selecting points from the original lidar point cloud $P_{d}$ that fall within the destination 3D bounding box. The edited range image $\mathbf{\tilde{x}}^{\text{(R)}}$ is resized to an object-centric range view using average pooling and denormalised before computing coordinate and intensity values (see Sec. B.2). We replace pixels in the original range view $R_{d}$ with the corresponding pixels from the edited range image if either (i) they fall within $\mathbf{m}_{\text{points}}$ or (ii) its corresponding 3D point in the edited range image is contained by the bounding box of the object, as seen in Fig. 5.
@@ 2.4 Training details
We consider objects from the nuScenes dataset [3] train split with at least 64 lidar points, whose 2D bounding box is at least $100\times 100$ pixels, with a 2D IoU overlap not exceeding 50% with other objects, and current camera visibility of at least 70%. Unless stated otherwise, our model is trained on “car” and “pedestrian” categories, dynamically sampling 4096 new actors per class each epoch. During training, once an object is selected, its current scene serves as the destination, from which we extract the 3D bounding box, environmental context, and ground truth insertion.
@@
Object references are taken from the same object at a different timestamp, picked randomly as follows. We collect references for the current object across all frames that meet the previous criteria to ensure good visibility and arrange them by normalised temporal distance $\Delta t$, where $1$ represents the furthest reference in time and 00 represents the current one. We then sample references randomly using a beta distribution $\Delta t\sim\text{Beta}(4,1)$ which ensures a preference for instances of the object that are far away from the current timestamp, see Sec. B for details.
@@ 3.1 Object insertion and replacement
To avoid situations where inpainted objects are placed at locations incompatible with the scene (e.g. a car on pavement), we use the position of existing objects and perform either object reinsertion or replacement, which differ by the choice of reference. This tests the model’s ability to generate realistic objects conditioned on a 3D bounding box while being semantically consistent with the scene. We sample 200 high-quality objects from the nuScenes val set as in Sec. 2.4, balanced across “car” and “pedestrian”.
@@
We define two types of references: same reference, where the source and destination images and bounding boxes are identical, meaning the object is reinserted in the same scene and position; and tracked reference, where the object is reinserted given its reference from a different timestamp, using the sampling strategy described in Sec. 2.4. This setting tests if the model can preserve the object’s appearance, and realistically perform novel view synthesis (for tracked reference).
@@
Results are presented in Fig. 3 both for replacement (rows 1–4) and insertion (row 5). We see that inpainted objects correspond tightly to their conditioning 3D bounding boxes while having a high degree of realism, both for camera (RGB) and lidar (depth and intensity), and show a strong coherence (lightning, weather conditions, occlusions, etc.) with the rest of the scene. The last row showcases object deletion, which can be achieved by using an empty reference image (note that we use empty references during training, as described in Sec. 2.4). Even though references in the replacement setting are from a different domain (time of day/weather), the model is able to inpaint such objects realistically. See Fig. S4 for more examples, including failure cases. We show an example of the composited camera and lidar scene in Fig. 5. We illustrate the flexibility of our bounding box conditioning and show that it is able to generate multiple views with a high degree of consistency, as illustrated in Fig. 4, Fig. S1 and Fig. S2.
@@
Figure 4: Our method can generate multiple novel views from a single reference image while maintaining multimodal consistency. From left to right: reference image $\mathbf{x}_{\text{ref}}$, extracted from a separate scene; original destination scene with the RGB image $\mathbf{x}^{\text{(C)}}$ and lidar range depth $\mathbf{x}_{0}^{\text{(R)}}$; and edited scenes. Note, the inpainted pedestrian moves to the right between frames, shifting the background to the left. Check Fig. S1 for extended results, including intensity.
Figure 5: Spatial compositing of camera-lidar object inpainting. Note that some background points are not overridden due to lidar reflections on the hood of the inserted car (bottom).
@@
Table 1: Adaptations of the pre-trained image VAE [21] from StableDiffusion [43] improving lidar reconstruction for depth (meters) and intensity (on a scale of $[0,255]$).
@@ Results
We report all realism metrics for camera-lidar object inpainting in Tab. 2 for the reinsertion and replacement settings. Compared to camera-only inpainting methods, MObI ($D=512$) achieves better results than PbE [65] across almost all benchmarks. We note that this method achieves competitive results in terms of FID, producing samples which are close in distribution to the target ones, yet LPIPS is much worse. This perceptual misalignment, which is more severe than even MObI (256) with no bounding box conditioning, could indicate that the use of joint generation of camera and lidar improves semantic consistency within the scene. We compare to simple copy&paste and show that this produces unrealistic composited images when replacing objects, even though it is sometimes used for training object detectors [12, 8, 52, 73]. Note that object reinsertion results for copy&paste, as well as CLIP-I scores, are not computed as the comparison would be unfair in this setting. We provide additional realism metrics in Tab. S1
We provide ablations for the 3D bounding box and the gated cross-attention adapter for $D=256$. Without the adapter, the box token is concatenated with the reference token in the PbE [65] cross-attention layer, followed by direct fine-tuning. Due to the lack of a baseline for lidar object inpainting realism, we provide comparative results for all experiments and ablation in the hope that this could constitute a baseline for future work. Comparing MObI (256) with both bbox conditioning and adapter to the one without bbox conditioning we notice significant improvements in perceptual alignment. Gated cross-attention leads to more realistic samples in the camera space, yet it does not improve lidar, hinting towards differences in training regimes for the two modalities. Finally, we note that realism scales strongly with resolution, leading us to believe that models operating at larger resolution would improve realism even more.
@@
Figure 6: Camera-lidar detection performance of an off-the-shelf BEVFusion [34] object detector on objects reinserted using our method. Left: we compute mAP at the scene-level, and TP errors (translation, scale, and orientation) on the reinserted objects only. Left: distribution of the scores of the true-positives show a modest shift towards lower scores for edited objects.
@@ 3.3 Object detection on reinserted objects
We compute mAP and error metrics on the re-inserted objects. Scene-level metrics such as mAP cannot be easily restricted to edited objects [^2] Because such metrics require false-positives which cannot be easily defined here. and will not be very sensitive to detection errors on these objects. We thus complement mAP with true-positive error metrics restricted to the re-inserted objects, computed following the usual matching procedure of the nuScenes devkit [3] but considering only ground-truth/detection pairs for inpainted objects.
@@ Results
Camera-lidar object detection results are presented in Fig. 6 (left), and we see that reinsertion comes at a small cost in object detection performance but that errors remain small (e.g. 0.161 AOE corresponds to a $9^{\circ}$ average error) while scene-level mAP is very similar. We also show the distribution of the scores of the true-positives in Fig. 6 (right), where the only scores suffer a modest decrease when the detector is applied to the reinserted samples. Overall, this shows that while a small domain gap exists, our method leads to samples that are both realistic and geometrically accurate, and that an off-the-shelf detector can successfully detect such objects even though it has not been trained on any synthetic data generated by our method. We show a sample of detections in Fig. S3 where the reinserted object is detected accurately and the bounding boxes of the untouched objects remain almost identical.
@@ 4 Strengths, limitations, and future work
Because we leverage a diffusion model pre-trained on webscale data, combined with the CLIP image encoding to guide the appearance of the inpainted object, our model can generate objects for classes unseen during training. In Fig. S5, we replace objects from nuScenes [3] classes beyond “car” and “pedestrian”, producing plausible, yet seemingly lower-quality results compared to seen classes.
Inpainting completely out-of-domain references produces unpredictable results. Since the model is finetuned on a narrow domain, the model reverts to known classes when faced with unfamiliar references, such as turning a horse into a brown car in Fig. S5 Extending our method to a true open-world setting by training on multiple 3D object detection datasets, as in [37], is an exciting future direction.
@@ A Extended Related Work
Synthetic lidar data generation has also advanced significantly. LidarGen [75] and LiDM [42] employ diffusion for lidar generation, with the latter also incorporating semantic maps, bounding boxes, and text. UltraLidar [63] densifies sparse lidar point clouds, while RangeLDM [19] accelerates lidar data generation by converting point clouds into range images using Hough sampling and enhancing reconstruction through a range-guided discriminator. DynamicCity [2] generates lidar occupancy grid sequences conditioned on dynamic scene layouts. However, full-scene generation can result in a large domain gap, particularly for downstream tasks like object detection, making it challenging to create realistic counterfactuals. [61, 22] focus on object-level lidar generation, with LOGen [22] using a point-based conditional diffusion framework, enabling fine-grained control over object geometry, viewpoint, and distance. None of the works mentioned above jointly generate camera and lidar data.
@@ B.3 Additional training details
Training takes approximately 20 hours on 8x 24GB NVIDIA A10G or 2x 80GB NVIDIA A100 GPUs. Inference throughput is about 8 camera+lidar samples per minute on a single A100.
@@
Ground Truth Vanilla Ours
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Original Reference Edited (C) Edited (R) - depth Edited (R) - intensity
Original Reference Insertion (C)
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Original Reference Replacement (C)
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption